It is reported that software developers need an average of 10 months to develop a software.



Assuming that the time taken by software developer is normally distributed with a standard



deviation of 2.3 months. Given that the probability that a developer needs to develop a



software more than c months is 0.985. Find the value of c.

1
Expert's answer
2021-12-16T15:37:16-0500

"P(Z>c)=0.985\to P(Z<c)=1-0.985=0.015."

"c=-2.17."

"\frac{x-10}{2.3}=-2.17."

"x=10-2.3*2.17=5" months.